Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report 2021-01-12

Observers:   Ken Smith, Michael Fulton
Support:  Shubham Srivastav, James Gillanders
TAT:  Sheng Yang, Cristina Barbarino and Janet Chen

EFOSC2 Calibrations:
Spectroscopic arcs
Biases 
SOFI Calibrations:
r-band morning flat

Weather:
Start: humidity 34% wind 0.6 m/s
Middle: humidity 14% wind 2.3 m/s
End: humidity 25% wind 3.5 m/s

Comments:



 UT Target Type Setup Slit Exposure[s] Airmass Seeing["] Notes
01:10Telescope problemsNot entirely sure what the problem was. Mention of "raptor".  So unable to start observations. I enquired whether switching to SOFI (backup plan) was possible, but we're informed that this would not make any difference. Looks like they had to reboot the telescope completely, so we lost about 1.5 hours.
02:39Telescope back up and running. Focus run at position of SN2021ek.
 02:50 SN2021ek Followup
 GR13 1 2700 1.3
 0.9
 
 03:41 SN2021ek Followup GR13 1 2700 1.5 0.9 
 04:29 GD71 STD GR11/GR13/GR16 1 4x100 1.6 0.8 
 04:46 SN2020znr Followup GR11 1 1800 1.7 1.2 
 05:20 SN2020znr Followup GR16 1 1800 1.7 1.1 
 06:02 AT2021wh Classification GR13 1600 1.0
 0.7SNIa, z=0.026, -7 to -4 days
 06:19SN2021dn  Followup GR11/GR16 1 1500/1500 1.8 0.7 
 07:33 FOCUS       
 07:43 SN2020ank Followup r-band dither  9 x 200 1.5 0.6PSF is not perfect, seeing very good. Ran out of time to fix, unfortunately.
08:22 PG1047 STD photometric field r-band  3 1.1 0.6 
 08:24 L745a STD GR11/GR13/GR16  4x100 1.5 0.6 
 08:46 AT2021zj Classification GR13  900 1.9 0.6SN IIn at -20 to -11 days. SNID z = 0.046, consistent with known host z = 0.0459.



Comments